OpenStudy (anonymous):

Match each equation with the description of the first step you can take to solve the equation. 1.) k/11 + 5/13 = -10/13 A.) subtract 5/13 from both sides 2.) k/11 - 10/13 = -5/13 b.) Add 5/13 to both sides 3.) -5/13 + k/11 = -10/13 c.) Add 10/13 to both sides 4.) 10/13 + k/11 = - 5/13 d.) subtract 10/13 from both sides

OpenStudy (texaschic101):

1. k/11 + 5/13 = - 10/13 (subtract 5/13 from both sides) 2. k/11 - 10/13 = - 5/13 (add 10/13 to both sides) 3. - 5/13 + k/11 = - 10/13 (add 5/13 to both sides) 4. 10/13 + k/11 = - 5/13 (subtract 10/13 from both sides)